The sum of Finn's age and Ellen's age is 40 years. Finn's age 10 years from now will be 1 year less than 4 times Ellen's age 6 years ago. Find their present ages.
Answer by venugopalramana(3286)

The sum of Finn's age (F)and Ellen's age(E) is 40 years.
F+E=40......................1
Finn's age 10 years from now
F+10
ELLENS AGE 6 YEARS AGO = E-6
4 TIMES THIS = 4(E-6)=4E-24
will be 1 year less than 4 times Ellen's age 6 years ago.
HENCE F+10=4E-24-1.....SUBSTITUTING FROM EQN.1
40-E+10=4E-25
5E=50+25=75
E=15
F=40-15=25

Find their present ages.